OpenStudy (anonymous):

ILL GIVE MEDAL TO WHOEVER SOLVES THIS! Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x − y = −11 2x + 3y = 5

rebeccaxhawaii (rebeccaxhawaii):

so you need the x and y intercepts ?

OpenStudy (anonymous):

everything you need is in the question

OpenStudy (jango_in_dtown):

where are the options? @Lord_Frieza

OpenStudy (ghostgate):

If you're trying to solve for 'x' and 'y' then I suppose you would use the elimination method to find them.

OpenStudy (ghostgate):

For instance you have these two equations: 4x − y = −11 2x + 3y = 5 Multiply the bottom equation by 2: 4x - y = -11 2(2x + 3y = 5) -> 4x + 6y = 10) Now subtract both equations: 4x - 4x = 0, -1y - 6y = -7y, -11 - 10 = -21 Now this is your equation: -7y = -21 y = 3 Now simply solve the other equation: 4x - 3 = -11 2x + 3(3) = 5 -> 2x + 9 = 5 [1] Add 3 to both sides. [2] Subtract 9 from both sides. 4x = -8 2x = -4 x = -2 So your conclusion: x = -2 and y = 3